OpenStudy (justuu):

Medal What is the elapsed time between 1:35 and 6:00? 5 hours and 35 minutes 4 hours and 35 minutes 5 hours and 25 minutes 4 hours and 25 minutes

OpenStudy (mathmate):

What happened to the crowd? No takers?

OpenStudy (mathmate):

Oh, it's not the same problem!

OpenStudy (mathmate):

In math, we don't guess! lol We work it out and compare which is the right answer! Do you know how to work it out?

OpenStudy (anonymous):

niether

OpenStudy (anonymous):

what's 1+5?

OpenStudy (mathmate):

What you do, is to write out the hours and minutes in two columns:

OpenStudy (justuu):

is it a right 6:00 - 1:35 =5:35

OpenStudy (mathmate):

h min 6 00 1 35 Since 00 cannot subtract 35, we borrow 1 hour and make it 60 minutes. h min 5 60 1 35 Now you do the subtraction, get the answer, and compare which choice. ok with that?

OpenStudy (mathmate):

Remember we borrowed one, so you have to subtract 1 hour from 5 hours. Look at the table above.

OpenStudy (anonymous):

the answer can't contain 5 hours because 1:35+5 hours=6:35. you've already gone over

OpenStudy (mathmate):

@KingOfNoOne I will let you continue, since you seem to be very interested in the task.... as long as you don't give a direct answer.

OpenStudy (anonymous):

It cant be 4 hours and 35 minutes either because 1:35+4:35=6:10